A motorboat travels 4km downstream in 0.5 h, and travels the same distance upstream in 2 h. Determine the speed of the boat in still water.
Answer by TimothyLamb(4379)

s = d / t
---
b + w = 4 / 0.5 = 8 kph
b - w = 4 / 2.0 = 2 kph
---
b + w = 8
b - w = 2
b = w + 2
---
b + w = 8
(w + 2) + w = 8
2w = 6
w = 3
---
Answer:
b = w + 2 = 5 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php